Status Report

NASA ISS On-Orbit Status 21 May 2008

By SpaceRef Editor
May 22, 2008
Filed under , , ,
NASA ISS On-Orbit Status 21 May 2008

All ISS systems continue to function nominally, except those noted previously or below.

Oleg Kononenko initiated the transfer of the potable water supplies brought up by Progress M-64 to the Service Module (SM)’s Rodnik water tanks (BV1, BV2). Later, the pumping equipment was dismantled and the activities were closed out. [After hooking up the plumbing connecting the 29P water tanks with the SM Rodnik tankage, the water was transferred at first in self-flow (under its own tank pressure), then using a compressor pump via a GZhS gas/liquid separator, to remove air bubbles in the water. The subsequent filling of the empty Progress tanks with urine will be scheduled later.]

After CDR Volkov prepared the auditory test equipment, he, FE-1 Kononenko & FE-2 Reisman took the periodic O-OHA (on-orbit hearing assessment) test, a 30-min. NASA environmental health systems examination to assess the efficacy of acoustic countermeasures, using a special MEC (Medical Equipment Computer) laptop application. It was the second session for the three crewmembers. [The O-OHA audiography test involves minimum audibility measurements for each ear over a wide range of frequencies (0.25-10 kHz) and sound pressure levels, with the crewmembers using individual-specific Prophonics earphones, Bose ANC headsets and the SLM (sound level meter). To conduct the testing, the experimenter is supported by special EarQ software on the MEC, featuring an up/down-arrow-operated slider for each test frequency that the crewmember moves to the lowest sound pressure level at which the tone can still be heard. The baseline test is required not later than about Flight Day 14 for each new Expedition and is then generally performed once per month. Note: There have been temporary hearing deficits documented on some U.S. and Russian crewmembers, all of which recovered to pre-mission levels.]

Reisman performed the periodic offloading of the Lab CCAA (Common Cabin Air Assembly) dehumidifier’s condensate tank, filling a CWC (Contingency Water Container) with the collected water slated for processing. No samples were required. [Transferred quantity is determined by allowing tank and CWC equalize with each other. Tank quantity stabilizes at a value higher than the neutral point.]

Kononenko serviced the Matryoshka-R (RBO-3-2) radiation payload, which has taken over the ESA/RSC-Energia experiment ALTCRISS (Alteino Long Term monitoring of Cosmic Rays on the ISS) with its AST Spectrometer, which Oleg rotated through 90 deg, and ALC (ALTCRISS laptop) equipment on DC1 panel 429.

The FE-2 conducted the periodic (monthly) CSA-CP (Compound Specific Analyzer-Combustion Products) maintenance/checkout, today on all four units. [The CSA-CP is a passive cabin atmosphere monitor that provides quick response capability during a combustion event (fire). Its collected data are stored on a logger. Garrett checked batteries, then zero-calibrated all instruments (to eliminate drift in the combustion sensors). Following zero calibration, the backup units were stowed in the Node, along with the sampling pump, while the prime unit was deployed at the SM Central Post.]

The CDR had an hour reserved for replacing and updating RODF (Russian Operations Data File) procedures pages.

Garrett relocated the IVCPDS/TEPC (Intravehicular Charged Particle Directional Spectrometer/Tissue Equivalent Proportional Counter), the primary radiation measurement tool in the ISS, from the COL (Columbus Orbital Laboratory) to the Lab at position LAB1S4, plugging its power cable in the CHeCS power/data outlet – port J1 on the PS-120 Junction Box at UOP #2. [TEPC was installed on the COL1A3 panel by Peggy Whitson at 3/3, powered from the J01 outlet on the COL1AD1 panel. Before that, Dan Tani had relocated TEPC to the SM (panel 410) on 1/23/08 after it had been running in Node-2 since 12/24/07.]

Sergey unpacked the newly arrived Sokol pressure suit from Progress M-64, conducted a leak check on it and then set it up in the Soyuz BO (Orbital Module) for airing out, i.e., drying out any humidity, later repeating the process on the Sokol gloves. Volkov also prepacked the Sokol with torn zipper for return to Earth. Afterwards, the equipment was stowed.

Reisman & Volkov concluded the Robotics ops begun yesterday by releasing the SSRMS LEE (Space Station Remote Manipulator System/Latching End Effector) at the Node-2 PDGF (Power & Data Grapple Fixture) and maneuvering it to the 1J Docking position. [The base change to the MBS (Mobile Base System) PDGF was completed yesterday by ground commanding.]

Oleg Kononenko had another 2 hrs for cargo transfers from 29P to the ISS, keeping track of the moves in the IMS (Inventory Management System).

The FE-2 performed the periodic deployment of four passive FMK (formaldehyde monitoring kit) sampling assemblies in the Lab (at P3, below CEVIS) and SM (at the most forward handrail, on panel 307) for two days, to catch any atmospheric formaldehyde on a collector substrate for subsequent analysis on the ground. [Two monitors each are usually attached side by side, preferably in an orientation with their faces perpendicular to the direction of air flow.]

In the US Airlock, Reisman continued EVA suit/tool battery maintenance, terminating the discharge process on the first batch of batteries and initiating it on the second group. [The charged batteries were removed from the BSA (Battery Stowage Assembly) and stowed in a bag.]

Later, Garrett had an hour set aside for his departure preparations.

As part of their standard fitness evaluation, Kononenko & Volkov undertook the Russian MO-5 MedOps protocol of Cardiovascular Evaluation during Graded Exercises on the VELO cycle ergometer, assisting each other as CMO (Crew Medical Officer). [The 50-min assessment, supported by ground specialist tagup via VHF and telemetry monitoring, uses the Gamma-1 ECG equipment with biomed harness, skin electrodes and a blood pressure and rheoplethysmograph cuff wired to the cycle ergometer’s instrumentation panels. For the graded exercise, the subject works the pedals after a prescribed program at load settings of 125, 150, and 175 watts for three minutes each. Data output involves a kinetocardiogram, rheoplethysmogram, rheoencephalogram and a temporal pulsogram.]

The CDR completed the routine maintenance of the SM’s SOZh/ECLSS system, including ASU toilet facilities systems/replaceables. [Regular daily SOZh maintenance consists of replacement of the KTO & KBO solid waste containers, replacement of an EDV-SV waste water and EDV-U urine container, and processing U.S. condensate water as it becomes available in a filled CWC from the Lab humidifier.]

The FE-1 performed the daily IMS (Inventory Management System) maintenance, updating/editing its standard “delta file” including stowage locations, for the regular weekly automated export/import to its three databases on the ground (Houston, Moscow, Baikonur).

The crew conducted their regular 2.5-hr. physical workout program (about half of which is used for setup & post-exercise personal hygiene) on the CEVIS cycle ergometer (FE-2), TVIS treadmill (CDR, FE-1), RED resistive exercise device (FE-2), and VELO bike with bungee cord load trainer (CDR/ MO-5, FE-1/MO-5).

Afterwards, Reisman transferred the crew’s exercise data file to the MEC for downlink, as well as the daily wristband HRM (Heart Rate Monitor) data of the workouts on RED, followed by their erasure on the HRM storage medium (done six times a week).

Hard-scheduled for Garrett today was the periodic audit/inventory of CDs in the CD Library II kit stowed in the Lab, going by a master list inside the case, to be hard-scheduled tomorrow of not completed today. [Any discrepancies and required updates were to be reported to MCC-H.]

Also on Reisman’s timeline today was the periodic survey/inspection of the contents of the CHeCS (Crew Health Care Systems) Rack.

A new item on the US “job jar” task list was the reconfiguration of food stowage containers in Node-2.

At ~12:55pm EDT, the FE-2 supported a live PAO TV interview with the Major League Baseball feature “This Week in Baseball” on Fox TV, to be aired before the “Saturday Game of the Week”.

Russian ASU Malfunction: While using the ASU toilet system in the SM, the crew heard a loud noise and the fan stopped working. After some troubleshooting the crew reported that the air/water Separator (MNR-RS) was not working. The crew then replaced the separator with a spare unit but reported afterwards that the ASU lacked suction. The crew next replaced the F-V filter insert, which provided good suction for a while but again exhibited weak suction. TsUP/Moscow instructed the crew to deactivate the ASU and use the toilet facility in the Soyuz spacecraft.

No CEO (Crew Earth Observations) photo targets uplinked for today.

CEO photography can be studied at this “Gateway” website: (as of 3/1/08, this database contained 757,605 views of the Earth from space, with 314,000 from the ISS alone).

Significant Events Ahead (all dates Eastern Time, some changes possible.):
05/31/08 — STS-124/Discovery/1J launch – JEM PM “Kibo”, racks, RMS (5:02pm EDT nominal)
06/02/08 — STS-124/Discovery/1J docking (1:49pm)
Tentative STS-124 docked working timeline:

  • 6/03 — EVA-1 (11:30am, 6.5 hrs.), OBSS transfer, JPM prep, S3/S4 SARJ TBA install, JPM install
  • 6/04 — JPM Vestibule outfitting, Reisman/Chamitoff handovers
  • 6/05 — EVA-2 (11:30am, 6.5 hrs), JTVE install, JRMS cvr remv, NTA prep, CP 9 ETVCG retrv
  • 6/06 — JLP relocate to JPM; JLP Vestibule leak check
  • 6/07 — JLP Vestibule outfitting
  • 6/08 — EVA-3 (10:30am, 6.3 hrs), S1 NTA install, CP9 ETVCG install, JRMS cover remove
  • 6/09 — JRMS checkouts, JLP Vestibule outfitting, A/L BCM R&R
  • 6/10 — Sayonara (~4:00pm), hatch close (~4:30pm)
  • 6/11 — Undocking (~9:04am); OBSS survey/inspection
  • 6/12 — Mostly off-duty
  • 6/13 — Stowing; deorbit preps
  • 6/14 — Deorbit burn (~9:56am);

06/14/08 — STS-124/Discovery landing (KSC: ~10:59am EDT, nominal)
07/10/08 — Russian EVA-20 (7/10-11)
??/??/08 — STS-125/Atlantis Hubble Space Telescope Service Mission 4 (SM4)
09/05/08 — ATV1 undocking
09/09/08 — Progress M-64/29P undocking (from DC1)
09/10/08 — Progress M-65/30P launch
09/12/08 — Progress M-65/30P docking
10/01/08 — NASA 50 Years
10/11/08 — Progress M-65/30P undocking (from SM aft port)
10/12/08 — Soyuz TMA-13/17S launch
10/14/08 — Soyuz TMA-13/17S docking (SM aft port)
10/16/08 — STS-126/Discovery/ULF2 launch – MPLM Leonardo, LMC
10/18/08 — STS-126/Discovery/ULF2 docking
10/23/08 — Soyuz TMA-12/16S undocking (FGB nadir)
11/03/08 — Soyuz TMA-13/17S relocation (from SM aft to FGB nadir)
11/20/08 — ISS 10 Years
11/26/08 — Progress M-66/31P launch
11/28/08 — Progress M-66/31P docking
12/04/08 — STS-119/Discovery/15A launch – S6 truss segment
12/06/08 — STS-119/Discovery/15A docking
12/15/08 — STS-119/Discovery/15A undocking
2QTR CY09 — STS-127/2J/A launch – JEM EF, ELM-ES, ICC-VLD
3QTR CY09 — STS-128/17A/Atlantis – MPLM(P), last crew rotation
05/??/09 — Six-person crew on ISS (following Soyuz 18S-2 docking)
3QTR CY09 — STS-129/ULF3/Discovery – ELC1, ELC2
4QTR CY09 — STS-130/20A/Endeavour – Node-3 + Cupola
1QTR CY10 — STS-131/19A/Atlantis – MPLM(P)
1QTR CY10 — STS-132/ULF4/Discovery – ICC-VLD, MRM1 (contingency)
2QTR CY10 — STS-133/ULF5/Endeavour – ELC3, ELC4 (contingency).

SpaceRef staff editor.